OpenStudy (anonymous):

-5t greater than or equal too 65

OpenStudy (anonymous):

help please!

OpenStudy (anonymous):

@pawtpie

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

\[-5t \ge 65\]

OpenStudy (anonymous):

If we want t on one side, what can we do?

OpenStudy (anonymous):

divide -5 by both sides and then u get x is greater than or equal too -13

OpenStudy (anonymous):

Very close

OpenStudy (anonymous):

When dealing with inequalities, when you multiply or divide by a negative number, the inequality sign flips.

OpenStudy (anonymous):

okay so what would it be then?

OpenStudy (anonymous):

\[-5t \ge 65\] Divided by 5: \[t \le-13\]

OpenStudy (anonymous):

Typo, that is divided by negative 5*